OpenStudy (anonymous):

tamara has two sisters. one of the sisters is 7 years older than tamara. the other sisters is 3 years younger than tamara. the product of tamara's sisters' age is 24. how old is tamara?

OpenStudy (mertsj):

Let tamara's age = x How do you represent her sister's ages?

OpenStudy (anonymous):

i put: t+7=x t-3=y xy=24

OpenStudy (mertsj):

Ok. Tamara's age is t The older sister is t+7 and the younger sister is t-3

OpenStudy (mertsj):

And if you multiply the older sister's age by the younger sister's age, you get 24.

OpenStudy (anonymous):

so if i do this: x-7=y+3 x-7=24/x+3 x+2-7x= 24+3x x+12 t+7=12 t+5 is tht right?

OpenStudy (mertsj):

There is no y. Where are you getting a y?

OpenStudy (mertsj):

(t+7)(t-3)=24

OpenStudy (anonymous):

i didnt get tht part thts why im asking

OpenStudy (mertsj):

Do you agree that the sister's ages are t t-7 and t-3 ?

OpenStudy (mertsj):

No. t+7

OpenStudy (anonymous):

yea

OpenStudy (mertsj):

The problem says to multiply her sisters' ages. Tamara is t years old. Her sisters are t+7 years old and t-3 years old. So if we are going to multiply her sister's ages, we must write (t+7)(t-3)

OpenStudy (anonymous):

oooh

OpenStudy (mertsj):

And the problem says that we will get 24 is we multiply those ages.

OpenStudy (mertsj):

\[t^2+4t-21=24\] \[t^2+4t-45=0\]

OpenStudy (mertsj):

Did you get it?

OpenStudy (p0sitr0n):

Tamara`s age is x. Her sisters age is hers + 7 so age+7 = x+7 her other sister is 3 years younger, so age - 3 = x-3 After they say the product is 24. so sister1 * sister2 = 24 (x+7) * (x-3) =24 Here we need to think. its a 2nd degree equation, so use quadratic formula

OpenStudy (p0sitr0n):

\[x^2+7x+3x+21 =24\] \[x^2 +10x -3 = 0\]

OpenStudy (anonymous):

The difference of the ages of the two sisters of tamara is 10. There product is 24 so the ages are 12 and 2. 12-2 =10 12x2 =24

OpenStudy (anonymous):

5

OpenStudy (anonymous):

let T = Tamara Sister 1 = S1 Sister 2 = S2 S1 = T + 7 S2 = T -3 ( T+7 )( T-3 ) = 24 T^2 + 4T - 45 = 0 ( T+9 )( T- 5 ) = 0 Discard T = - 9 Tamara = 5 S1 = 5+ 7 = 12 S2 = 5 - 3 = 2

OpenStudy (anonymous):

@Luis_Rivera you do realy need to compute the age of Tamara

OpenStudy (anonymous):

You do not

OpenStudy (anonymous):

pretty, isnt it ?